OpenStudy (anonymous):

solve the sytem by substitution x-2y=2 2x=3y=11 solve the sytem by substitution x-2y=2 2x=3y=11 @Mathematics

OpenStudy (anonymous):

are you sure this equation about 2x=3y=11 ?

OpenStudy (anonymous):

i corrected it sorry it is x-2y=2 2x+3y=11

OpenStudy (anonymous):

x=2y+2 2(2y+2)+3y=11 -> 4y+4+3y=11 7y=7 -> y=1 then find x x=2(1)+2 -> x=4

OpenStudy (anonymous):

ok so how do you do it if it is like this. y=3x x+y=4

OpenStudy (anonymous):

substitute to x+y=4 x+(3x)=4 -> 4x=4 x=1

OpenStudy (anonymous):

ok, i kind of understand what your doing, it is the rule of the substition i am kind of getting confused about. i see how you eliminate and then divide at the end right? but exactly how are you combining the sets?

OpenStudy (anonymous):

just replace the variable

OpenStudy (anonymous):

or what about a fraction? y= -x/2 2x-3y= -7

OpenStudy (anonymous):

it mean replace y with -x/2 then substitution to 2x-3y=-7

OpenStudy (anonymous):

i am having a hard time cuz i was out sick with bronchitis this last week. then the teacher says , here is the homework. get it done so now i am lost.

OpenStudy (anonymous):

change y for the second equation with -x/2 so 2x-3(-x/2)=-7 2x+3x/2=-7 4x/2 +3x/2 =-7 7x/2=-7 7x=-14 -> x=-2